SerCP10 1P015 fathom is a unit of length, usually reserved for measuring the depth of water. A fathom is approximately o ft in length. Take the distance from Earth to the Moon to be 246,000 miles nd use the given approximation to find the distance in fathoms. (Enter a number with the Your answer cannot be understood or graded. More Information fathoms
Answer #1

1 fathom = 6ft

Distance between earth and moon, d = 246,000 miles

1mile = 5280 ft = (5280/6) fathom = 880 fathom

So, 1mile = 880 fathom

Hence, distance between earth and moon = 880*246000 fathoms = 216,480,000 fathoms